OpenStudy (raphaelll):

A ball is dropped from a building. The distance, D metres, that it falls in t seconds is given by D=5tsquared Find the distance fallen in 3 seconds Find the time taken to fall 100 metres (answer to 1 decimal place)

OpenStudy (danjs):

given: \[Distance = 5*(time)^2\] that right?

OpenStudy (danjs):

\[D(t) = 5*t^2\]

OpenStudy (danjs):

For the first part, uses time = 3 and solve for D For the second, use D=100, solve for time t

OpenStudy (raphaelll):

No, D = 5t squared

OpenStudy (danjs):

that is what i typed \[D = 5t^2\] "D euals 5 tee squared"

OpenStudy (danjs):

TO find the distance fallen in 3 seconds, put t=3, solve for D=5*3^2 to find the time to fall 100 meters put, 100 = 5*t^2, solve for t

OpenStudy (raphaelll):

ah i see thank you i've got it